A compound with a molar mass of 544.0 g/mol is made up of 26.5 grams carbon, 2.94 grams Hydrogen and 70.6 grams Oxygen. What is its empirical and molecular formula?

1
Expert's answer
2021-11-22T05:59:01-0500

Molar mass of Carbon = 12.011

26.5/12.011=2.21mol

Molar mass of H=1.00784

2.94/1.00784=2.92

Molar mass of O= 15.99

70.6/15.99= 4.42

CHO2